# Standard Gaussian Process Can Be Excellent for High-Dimensional Bayesian Optimization

### v3

@@ 3.2 Structural Assumption in Input Space
Low-dimensional Embedding. Wang et al., [32] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [25] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
Partial Variable Dependency. Another strategy is to trim down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [11]. While one can also use MAP estimation, the performance is often much worse.
@@ 4 Comprehensive Evaluation
We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic experiments.
@@ 4.1 Experimental Settings
We employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
@@
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section A.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 2.
Real-World Benchmarks. We employed the following real-world benchmark problems.
@@
  • Mopta08 (124)  [13], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed [6] to encode the constraints as a soft penalty, which is added into the objective.
  • SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
  • Rover (60), a rover trajectory planning problem from [31]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
  • DNA (180) [35], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [22]. Prior analysis [35] shows that only around 43 variables are relevant to the target function.
  • NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
Method Structural Assumption
ALEBO, HESBO Low-dim embedding
SaasBO Partial variable dependency
RDUCB, Tree UCB Additive function decomposition
SBO, TURBO, BNN None
Methods. To execute the standard BO (SBO), we used GP regression with ARD Matérn-5/2 kernel (see (1)). We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch [9] for surrogate training. To compare with other high-dimensional BO methods, we considered two training schemes. One performs MAP estimation with LBFGS. The maximum number of iterations was 15K and tolerance level 2E-9. We employed a diffuse prior over each length-scale, i.e., Uniform(0.001, 30). The noise variance was assigned a diffuse Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We denote this method as SBO-1. The second training scheme simply uses ADAM optimization [17] to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We denote this method as SBO-2. We also tested with Matérn-5/2 kernel with a single length-scale parameter, which we denoted as SBO-3. We used BOTorch [2] for acquisition function maximization, which uses LBFGS. Additional evaluations on alternative choices of the kernel, training scheme, and acquisition function are provided in Appendix Section B.
@@
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [34], HESBO [25], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. We used the original implementation and default (recommended) settings for the competing methods. In addition, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Table 3 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section A.2.
@@
For each optimization task, we randomly queried the target function for 20 times to generate the initial data. We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.

### v4

@@ 3 High Dimensional Bayesian Optimization
Structural Assumption in Input Space. Another class of methods assumes a low-rank structure in the input space. Many of them introduce Low-dimensional Embeddings. Wang et al., [31] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [23] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue. Another direction is to impose Partial Variable Dependency by triming down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [12]. While one can also use MAP estimation, the performance is often much worse.
4 Theoretical Analysis
To understand why standard BO can fail in high-dimensional cases, in particular with the SE kernel (see Section 6), we look into the gradient structure during the GP training. Specifically, we consider at the beginning, all the length-scale parameters are set to the same initial value, $\ell_{1}=\ldots=\ell_{d}=\ell_{0}$. From the marginal likelihood (3), we can derive the gradient w.r.t each length-scale,
$\frac{\partial\log p({\bf y}|{\bf X})}{\partial\ell_{k}}=\frac{1}{2}{\rm tr}\left({\bf A}\cdot\frac{\partial{\bf K}}{\partial\ell_{k}}\right),$ (5)
where ${\bf A}=\bm{\alpha}\bm{\alpha}^{\top}-({\bf K}+\sigma^{2}{\bf I})^{-1}$, and $\bm{\alpha}=({\bf K}+\sigma^{2}{\bf I})^{-1}({\bf y}-{\bf m})$. When using the SE kernel as specified in (1), each entry of $\frac{\partial{\bf K}}{\partial\ell_{k}}$ is as follows: $\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ii}=0$ and for $i\neq j$,
$\left[\frac{\partial{\bf K}}{\partial\ell_{k}}\right]_{ij}=\frac{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})}{\partial\ell_{k}}=\frac{2a}{\ell_{k}}\exp(-\frac{\|{\bf x}_{i}-{\bf x}_{j}\|^{2}}{\ell_{0}^{2}})\frac{(x_{ik}-x_{jk})^{2}}{\ell_{k}^{2}}\leq\frac{2a}{\ell_{k}}\cdot\frac{\rho^{2}}{e^{\rho^{2}}},$ (6)
where $x_{ik}$ and $x_{jk}$ are the $k$-th element in ${\bf x}_{i}$ and ${\bf x}_{j}$, respectively, and $\rho=\frac{\|{\bf x}_{i}-{\bf x}_{j}\|}{\ell_{0}}$.
Intuitively, as the input dimension $d$ increases, the squared distance between input vectors grows rapidly. Consequently, the factor $\frac{\rho^{2}}{e^{\rho^{2}}}$ in (6) can quickly fall below the machine epsilon $\xi$, causing the gradient vanishing issue [3, 11]. That is, the length-scale cannot be effectively updated according to $\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}$, due to the limitations of the numerical precision.
Proposition 4.1.
Given any $\xi>0$, $\frac{\rho^{2}}{e^{\rho^{2}}}<\xi$ when $\rho>\tau_{\text{SE}}=\frac{1}{2}+\sqrt{\frac{1}{4}-\log\xi}$.
When using the double-precision floating-point format (float64), the rounding machine epsilon is $\xi=2^{-53}$, and the above threshold is $\tau_{\text{SE}}=6.58$. To analyze the likelihood of gradient vanishing, we provide a probabilistic bound under uniform data distribution.
Lemma 4.2.
Suppose the input domain is $[0,1]^{d}$ and the input vectors are sampled independently from the uniform distribution, then for any constant threshold $\tau>0$, when $d>6\ell_{0}^{2}\tau^{2}$,
$p(\rho\geq\tau)>1-2\exp\left(-\frac{(d-6\ell_{0}^{2}\tau^{2})^{2}}{18d}\right).$ (7)
The lower bound grows exponentially with the input dimension $d$ and converges to one. This implies that given any fixed choice of $\ell_{0}$ relevant to $d$ (e.g., $\ell_{0}=1$), as $d$ increases, the probability $p(\rho\geq\tau_{\text{SE}})$ will rapidly approach one, leading to vanishing gradient for every $\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})$. As a consequence, each length scale $\ell_{k}$ in (5) cannot be effectively updated, preventing meaningful training [^1] A union bound can be directly derived from (7) to lower bound the joint probability that all gradients $\{\partial\kappa_{\text{SE}}({\bf x}_{i},{\bf x}_{j})/\partial\ell_{k}\}_{i,j}$ across $N$ training instances are below machine epsilon. However, this bound is too loose to reflect our practical observation (see Fig. 1 and  2 and Table 2). Therefore, we focus our discussion on a single pair of inputs. . The inferior prediction accuracy of the GP model further contributes to poor performance of the BO procedure. Since $\tau_{\text{SE}}$ is small, gradient vanishing can occur even when $d$ is not large. For example, for $\ell_{0}=0.5$ and $\ell_{0}=1$, we have $p(\rho>\tau_{\text{SE}})>0.99$ when $d\geq 205$ and $d\geq 473$, respectively.
Why Matérn kernels perform much better? In our practical evaluations (see Section 6), standard BO with Matérn kernels (2) often perform much better than with the SE kernel in high-dimensional problems, frequently achieving state-of-the-art results. To understand the reasons behind this improvement, we apply the same framework to analyze the behavior of GP training with the Matérn kernel. First, we obtain the gradient,
[…]
As the distance increases, the factor $\rho^{2}/e^{\sqrt{5}\rho}$ will also converge to zero, implying that with the growth of $d$, the gradient will eventually fall below machine epsilon. However, since this factor $\rho^{2}/e^{\sqrt{5}\rho}$ decreases much slower than the corresponding factor to the SE kernel, namely $\rho^{2}/e^{\rho^{2}}$ in (6), the Matérn kernel is able to robustly handle much higher dimensions.
Proposition 4.3.
[…]
Lower bound 0.95 0.99 0.995 0.999 0.9995 0.9999
SE ($d$) 172 205 219 250 264 294
Matérn ($d$) 980 1040 1064 1116 1137 1185
Table 1: The input dimension $d$ vs. lower bound probability of gradient vanishing under $\ell_{0}=0.5$.
Numerical Verification. To verify whether gradient vanishing occurs during practical GP training and to access whether our analysis aligns with the empirical observations, we conducted numerical experiments using two synthetic functions, Hartmann6($d$, 6) and Rosenbrock(d,d) (see Section 6.1 for definitions). We varied the input dimension $d$ across {50, 100, 200, 300, 400, 500, 600}, with the domain set to $[0,1]^{d}$. For each value of $d$, we uniformly sampled $500$ training inputs and $100$ test inputs. We trained GP with both SE and Matérn kernels, using length-scale initialization $\ell_{0}$ from $\{0.1,0.5,0.693,1.0,\sqrt{d}\}$. Note $\ell_{0}=0.693=\text{SoftPlus}(0)$ is a popular initialization choice, used as the default choice in influential GP/BO libraries GPyTorch [9] and BOTorch [2]. Training and testing were repeated 20 times for each $d$ and $\ell_{0}$. We then evaluated (1) the average relative $L_{2}$ difference between the length-scale vector before and after training, namely, $\|\bm{\ell}_{\text{trained}}-\bm{\ell}_{\text{init}}\|/\|\bm{\ell}_{\text{init}}\|$, (2) the average gradient norm at the first training step, and (3) the average test Mean-Squared-Error (MSE).
As shown in Fig. 1, gradient vanishing indeed occurred, leading to training failure. When using $\ell_{0}=0.1$ for GP training with the SE and Matérn kernels, both the length-scale relative $L_{2}$ difference and the gradient norm remained close to zero across all choices of $d$ (those values are not displayed because they are too small). Consequently, the MSE hovered around 1.0, indicating consistent training failure. When we increased $\ell_{0}$ to 0.5 and 1.0 for training with the SE kernel, both the length-scale difference and gradient norm became significantly larger and non-trivial for $d=50$ and $d=100$, resulting in MSE below 0.2. This demonstrates that gradient vanishing did not occur and training was successful. However, as $d$ increased, gradient vanishing reappeared at $d=200$ for $\ell_{0}=0.5$ and at $d=400$ for $\ell_{0}=1.0$, for both Hartmann6 and Rosenbrock. In these cases, the length-scale difference and gradient norm dropped near zero, while the MSE returned to around 1.0. A similar trend was observed with the Matérn kernel. However, GP training with the Matérn kernel can succeed with higher dimensions under the same initialization conditions. For example, with $\ell_{0}=0.5$, for both Hartmann6 and Rosenbrock, training with the Matérn kernel started to failed until $d=400$, compared to $d=200$ with the SE kernel. When using $\ell_{0}=1.0$, the Matérn kernel succeeded across all $d$ values, while the SE kernel started to fail at $d=400$. Table 2 summarizes the input dimensions at which gradient vanishing and training failure began to happen.
Confirmation in BO Running. We next investigated whether gradient vanishing occurs during the BO process, potentially leading to poor optimization performance. To this end, we examined two benchmark BO tasks: SVM (D=388) and Rosenbrock (300,100), as detailed in Section 6. As shown in Fig. 3 and 4, standard BO with the Matérn kernel (SBO-Matérn) achieved top-tier results, while BO with the SE kernel (SBO-SE) performed poorly. Both methods used the initialization $\ell_{0}=0.693$. We then extracted the BO trajectories for both methods and analyzed the relative $L_{2}$ difference in the length-scale vectors before and after training, and the gradient norm at the first training step. As illustrated in Fig. 2, during the BO iterations, GP-SE ($\ell_{0}=0.693$) consistently experienced gradient vanishing and so training failure, whereas GP-Matérn ($\ell_{0}=0.693$) maintained non-trivial gradients. This difference directly correlates with their respective optimization performances — poor for SBO-SE and excellent for SBO-Matérn.
Figure 1: Mean Square Error (MSE), Relative $L_{2}$ difference between the length-scale vectors before and after training, and $L_{2}$ norm of the length-scale gradient at the first training step, across different input dimensions ($x$-axis) and initializations (legend). The results were averaged from 20 runs. For $\ell_{0}=0.1$, the length-scale relative difference and gradient norm are not displayed, as they fall much below the minimum values on the $y$-axis.
Figure 2: The relative $L_{2}$ difference of the length-scale vectors before and after training and the gradient norm at the first training iteration across every step of BO. The results for GP-SE ($\ell_{0}=0.693$) are not displayed, as they fall much below the minimum values on the y-axis.
Hartmann6 $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
SE 50 200 200 400
Matérn 50 400 600
Rosenbrock $\ell_{0}=0.1$ 0.5 0.6931 1.0 $\sqrt{d}$
SE 50 200 200 400
Matérn 50 400 600
Table 2: The input dimension from which gradient vanishing started to occur and training started to fail; ✓means no gradient vanishing occurred.
5 Robust Initialization
From the analysis in Section 4, one straightforward approach to improving learning with the SE (and also Matérn) kernel is to initialize each length-scale with a larger $\ell_{0}$, which can loosen the lower bound in (7) and allow the GP model to better adapt to higher dimensions. However, this approach is not robust enough, as $d-6\ell_{0}^{2}\tau^{2}$ will continue to increase with $d$, and the bound will still rise toward one rapidly, causing the gradient to vanish. To address this issue, we propose setting $\ell_{0}$ to growing with $d$ rather than keeping it constant,
$\ell_{0}=c\sqrt{d},\;\;\;c>0.$ (8)
Lemma 5.1.
Suppose the input domain is $[0,1]^{d}$ and each input vector is independently sampled from the uniform distribution. Given any constant threshold $\tau>0$, we set $\ell_{0}=c\sqrt{d}$ such that $c>\frac{1}{\sqrt{6}\tau}$, then
$p(\rho\geq\tau)\leq 2\exp\left(-2(c^{2}\tau^{2}-\frac{1}{6})^{2}d\right)\propto\exp(-\mathcal{O}(d)).$ (9)
With this initialization, the increase in $d$ exponentially reduces the upper bound on the probability of $\rho$ exceeding the given threshold $\tau$. As a result, the gradient vanishing issue can be fundamentally mitigated in the high-dimensional settings.
In our experiments, we tested standard BO using SE kernel with our proposed length-scale initialization as specified in (8) where we set $c=1$. As shown in Fig. 3 and Fig. 4, the performance across all the benchmarks is dramatically improved, as compared to using popular initialization $\ell_{0}=0.693$. The performance matches that of using Matérn kernels, achieving state-of-the-art (or near-best) results. In addition, as shown in Fig. 1 and Table 2, our numerical experiments confirm that gradient vanishing never occurred with our robust initialization method. Fig. 2 further demonstrates that our method eliminates gradient vanishing during practical BO runs, enabling effective GP training and much better optimization results. Together this shows that our method can effectively adapt SE kernels to high-dimensional settings, and mitigate the failure mode of standard BO.
Alternative method. Recent work [13] proposes an intuitive approach to address the learning challenge of the SE kernel in the high-dimensional spaces by constructing a log-normal prior over each length scale, $p(\ell_{k})=\text{LogNormal}(\ell_{k}|\mu_{0}+\frac{\log(d)}{2},\sigma_{0})$. This prior regularizes the length-scale to be on the order of $\sqrt{d}$ during training. Using our theoretical framework, this method can be justified as an alternative strategy to alleviate the gradient vanishing issue. However, our method, directly motivated by our analysis, is even simpler — requiring no additional prior construction or associated parameters, and has also shown effective in addressing the gradient vanishing issue. The comparison with  [13] is given in Section 6.
@@ 6.1 Experimental Settings
For a comprehensive evaluatio of the standard BO, we employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity, all the tasks aim for function maximization.
@@
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section B.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 3.
@@
Real-World Benchmarks. We employed the following real-world benchmark problems: Mopta08 (124)  [14], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints.
We followed [6] to encode the constraints as a soft penalty, which is added into the objective. SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model. Rover (60), a rover trajectory planning problem from [30]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment. DNA (180) [26], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [20]. Prior analysis [26] shows that only around 43 variables are relevant to the target function. NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
Methods. We tested with three versions of standard BO:
  • SBO-Matérn: GP regression with ARD Matérn-5/2 kernel as specified in (2), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$, as the default choice in the influential GP/BO libraries GPyTorch and BOTorch.
  • SBO-SE: GP regression with ARD SE kernel as specified in (1), the length-scale initialization $\ell_{0}=\text{SoftPlus}(0)=0.693$.
  • SBO-SE(RI): GP regression with ARD SE kernel, using our proposed robust initialization $\ell_{0}=c\sqrt{d}$ where $d$ is the input dimension and $c=1$
We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch for GP training. To expose the possible gradient vanishing issue , we used Uniform(0.001, 30) over each length-scale. The noise variance was assigned a diffused Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We ran L-BFGS to obtain MAP estimation. The maximum number of iterations was 15K and tolerance level 2E-9.
@@
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [33], HESBO [23], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. In addition, we tested the most recent method VBO (Vanilla BO[13] that levels up the performance of the standard BO by constructing a log-normal prior over the length-scale parameters (see Section 5). We used the original implementation and default (recommended) settings for the competing methods. Besides, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Appendix Table 6 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section B.2.
@@
For each optimization task, we randomly queried the target function for 20 times to collect the initial data. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.